### v1

@@ 3.3 Training Pipeline
The overall training pipeline of our Croc model is shown in Fig. 2. Building upon LLaVA-1.5, the Croc undergoes a two-stage pre-training procedure followed by instruction tuning.
@@
Stage 1.5: Cross-Modal Comprehension. Building on the cross-modal alignment stage, we introduce the cross-modal comprehension phase as a subsequent pre-training stage. To facilitate comprehensive learning of visual tokens by the LLM, we pretrain the projection layer as well as the LLM in this stage. We select 1.2M detailed image-text pairs from the ShareGPT4V (Chen et al., 2024b) dataset, as the detailed captions (as shown in Fig. 4) can enhance the visual token learning of the LLM. To prevent the degradation of the inherent capabilities of LLM, we also incorporate 300K pure text data. Please refer to the appendix for more details.
@@ 4 Experiments
| Method | LLM | Res. | Pretrain | Finetune | VQAv2 | GQA | VizWiz | SciQA-I | TextVQA |
| BLIP-2 | Vicuna-13B | 2242 | 129M | – | 65.0 | 41.0 | 19.6 | 61.0 | 42.5 |
| InstructBLIP | Vicuna-7B | 2242 | 129M | 1.2M | – | 49.2 | 34.5 | 60.5 | 50.1 |
| InstructBLIP | Vicuna-13B | 2242 | 129M | 1.2M | – | 49.5 | 33.4 | 63.1 | 50.7 |
| Shikra | Vicuna-13B | 2242 | 600K | 5.5M | 77.4∗ | – | – | – | – |
| IDEFICS-9B | LLaMA-7B | 2242 | 353M | 1M | 50.9 | 38.4 | 35.5 | – | 25.9 |
| IDEFICS-80B | LLaMA-65B | 2242 | 353M | 1M | 60.0 | 45.2 | 36.0 | – | 30.9 |
| Qwen-VL | Qwen-7B | 4482 | 1.4B† | 50M | 78.8∗ | 59.3∗ | 35.2 | 67.1 | 63.8 |
| Qwen-VL-Chat | Qwen-7B | 4482 | 1.4B∗ | 50M | 78.2∗ | 57.5∗ | 38.9 | 68.2 | |
| LLaVA-1.5-7B | Vicuna-7B | 3362 | 558K | 665K | 78.5∗ | 62.0∗ | 50.0 | 66.8 | 58.2 |
| LLaVA-1.5-13B | Vicuna-13B | 3362 | 558K | 665K | 80.0∗ | 63.3∗ | 61.3 | ||
| Croc-7B | Vicuna-7B | 3362 | 558K+1.5M | 665K | 64.2∗ | 50.0 | 70.1 | 60.4 | |
| Croc-13B | Vicuna-13B | 3362 | 558K+1.5M | 665K | 80.7∗ | 57.1 | 72.7 | 60.8 |
@@
| BLIP-2 | Vicuna-13B | 2242 | 129M | – | – | – | – | – |
| InstructBLIP | Vicuna-7B | 2242 | 129M | 1.2M | – | – | – | – |
| InstructBLIP | Vicuna-13B | 2242 | 129M | 1.2M | – | – | – | – |
| Shikra | Vicuna-13B | 2242 | 600K | 5.5M | – | – | – | – |
@@ 4.1 Implementation Details and Evaluation Benchmarks
Following LLaVA-1.5, we utilize CLIP ViT-L/14@336px (Radford et al., 2021) as the visual encoder and the Vicuna (Chiang et al., 2023) 7b/13B model as the LLM. During the cross-modal alignment stage, the learning rate for the projection layer is set to . In the cross-modal comprehension stage, the learning rates are adjusted to for both the LLM and projection layer and for the prompt token pool and prediction layer. The random mask ratio is set as and the size of the prompt token pool is set to . In the instruction tuning stage, we adopt the learning rate of for LLM and the projection layer. Following previous works (Yang et al., 2023a; An et al., 2023; Gu et al., 2024; Yang et al., 2024), we employ AdamW (Loshchilov & Hutter, 2019) as the optimizer, initialized with a weight decay of . The parameters and are set to and , respectively. We train Croc on NVIDIA A100 (80G) GPUs.
@@
To prove the effectiveness of the Croc model, we evaluate our models across various benchmarks, including 1) General Visual Question Answering: GQA (Hudson & Manning, 2019), VQAv2 (Goyal et al., 2017b), VizWiz (Gurari et al., 2018), LLaVA-Bench (In-the-Wild) (Liu et al., 2024b), MM-Vet (Yu et al., 2024), RealWorldQA; 2) OCR-Related Question Answering: TextVQA (Singh et al., 2019), DocVQA (Mathew et al., 2021); 3) Illusion Benchmarks: POPE (Li et al., 2023c); 4) Comprehensive Reasoning Benchmarks: MMBench (Liu et al., 2024c), MME (Yin et al., 2023), MMStar (Chen et al., 2024a), MM-Vet (Yu et al., 2024); 5) Science Visual Question Answering: SciQA (Lu et al., 2022), AI2D (Kembhavi et al., 2016);
4.2 Quantitative Analysis
@@
In Tab. 1, we present a detailed comparison between our model and previous state-of-the-art methods across various benchmarks tailored for academic tasks. Croc outperforms existing models on 4/5 benchmarks. Specifically, compared to LLaVA-1.5-7B, Croc-7B demonstrates performance improvements of 2.0%, 2.2%, 3.3%, and 2.2% on VQAv2, GQA, SciQA-I, and TextVQA. Similarly, Croc-13B shows improvements of 0.7%, 0.7%, 3.5%, and 1.1% over LLaVA-1.5-13B on VQAv2, GQA, VisWiz, and SciQA-I. Additionally, we compare Croc with baselines on instruction-following benchmarks. As shown in Tab. 2, Croc achieves superior performance on 6/7 datasets. Notably, compared to LLaVA-1.5-7B, Croc-7B exhibits significant performance gains of 1.0%, 15.7, 3.3%, 1.4%, 5.5%, 8.5%, and 4.4% on POPE, MME, MMB, MMB-CN, SEED, LLaVA-W, and MM-Vet. Furthermore, we evaluate Croc’s performance on the multimodal understanding benchmarks in Tab 3. Croc-7B obtains significant improvements of 7.4%, 4.4%, 3.8%, and 1.5% over LLaVA-1.5-7B on MMStar, DocVQA, AI2D, and RealWorldQA, respectively. These performance improvements are primarily attributed to the proposed novel cross-modal comprehension pretraining stage, which significantly enhances the LLM’s capability to integrate and understand visual tokens.
4.3 Qualitative Analysis
@@
Attributable to the cross-modal comprehension stage, our proposed Croc model adeptly captures fine-grained semantic information from images, significantly enhancing its visual understanding and reasoning capabilities. In the first example shown in Fig. 5, unlike LLaVA-1.5, our model captures more detailed information from images, such as specific vehicle models like Audi and lower shooting angles. As illustrated in the second example of Fig. 5, our Croc model efficiently identifies and describes detailed semantic information, such as the “parachute” under the aircraft.
@@ 4.4 Ablation Study
Prompt Token Pool Size. The prompt token pool is designed to replace part of the original visual tokens with the most relevant prompt tokens. As shown in Fig. 6(a), the larger the number of tokens in the pool, the easier it is to generate detail captions, resulting in a lower training loss (). The decrease in training loss is saturated when the pool size is 2,048. When the pool size is 1, 75% of the image tokens are directly dropped and the training is too difficult, resulting in poor performance in Tab. LABEL:table:pool_size. We observe that increasing the token pool from 1 to 2,048 tokens improves performance. However, expanding the pool further to 4,096 tokens degrades performance because the visual token reconstruction task is too easy to improve the detailed caption generation task.
Visual Token Mask Ratio. The mask ratio of visual tokens directly affects the difficulty of the visual token reconstruction task, and thus significantly affects the effectiveness of pretraining. In Table. LABEL:table:mask_ratio, we report the results of experiments with different mask ratios. Similar to the observations with MAE (He et al., 2022), we find that using a 75% masking ratio yields optimal results in several downstream benchmarks. Lower ratios, e.g., 0%, make the pretraining task too easy, while higher ratios, e.g., 90%, make the reconstruction task too difficult. Both extremes lead to reduced pretraining effectiveness.
Nearest Neighbors v.s. Hungarian Matching. To improve the utilization rate of the tokens in the prompt token pool, we utilize Hungarian Matching to replace 75% of visual tokens with the most relevant prompt tokens. In Fig. 6(b), we present a comparative analysis of token utilization using both Nearest Neighbors and Hungarian Matching. Due to the Hungarian Matching algorithm’s stringent requirement to select distinct prompt tokens, there has been a significant improvement in the overall utilization of prompt tokens. This enhancement of utilization substantially improves the representational capabilities of the model (Zhu et al., 2024). Therefore, as shown in Tab. LABEL:table:matching, employing the Hungarian Matching algorithm, in contrast to the Nearest Neighbors algorithm, yields substantial performance enhancements across all evaluated benchmarks.
Unidirectional Attention v.s. Mixed Attention. To verify the impact of the mixed attention mechanism, we conduct experiments to compare mixed attention with unidirectional attention used in LLaVA-1.5. As shown in Tab. LABEL:table:attention, our proposed mixed attention mechanism achieves significant performance improvement on all the benchmarks. To further explore the influence of different attention mechanisms, we compare the average attention scores of response tokens to visual tokens and language instruction tokens in Fig. 7. Here, the same question “Please describe in detail what is in the picture.” is used for 200 randomly selected images. The Croc-7B model exhibits significantly higher attention scores for visual tokens compared to the LLaVA-1.5-7B model, as indicated by the brighter region in the visual token part of the heatmap. This suggests that Croc-7B puts more emphasis on understanding and attending to the image content when generating responses.
Brief Caption v.s. Detailed Caption. To investigate the impact of detailed captions on our proposed cross-modal comprehension stage, we employ the BLIP model to generate brief captions for the 1.2M images from the ShareGPT4V dataset. As shown in Tab. LABEL:table:caption, the use of detailed captions achieves a significant performance improvement. This is mainly because the detailed caption contains rich semantic details of images (as shown in Fig. 4), facilitating visual token reconstruction by LLM.
Ablation on Stage, Data, and Objective. In Tab. LABEL:table:stage1, we present the performance results without the cross-modal alignment pretraining stage. The misalignment between the features of the LLM and the visual encoder significantly increases the difficulty of the visual token reconstruction task. Consequently, the performance of the model on all the benchmarks is reduced considerably.
In Tab. 5, we present a series of comprehensive ablation studies to elucidate the influence of pre-training data and different training objectives. Initially, we augment the 558K pertaining data of LLaVA-1.5 with the additional 1.5M data and apply the same training method as LLaVA-1.5. This augmentation leads to performance improvements in only 4/9 datasets, while the others exhibit declines. This phenomenon can be attributed to the significant alteration in the distribution of the original 558K pretraining dataset caused by the introduction of the additional data. Building upon this foundation, we incorporated a visual token reconstruction task, which facilitated the establishment of more effective integration between the LLM and visual tokens. This integration led to improved performance across 5/9 datasets. Additionally, benefiting from the detailed description captions, we observe performance improvements in 8/9 datasets after implementing the detailed caption generation task. Finally, by combining visual token reconstruction and detailed caption generation tasks, the detailed caption further enriches the LLM’s understanding of visual features, resulting in significant performance enhancements in our model.
@@
VILA v.s. Croc. As VILA incorporates an additional one million instruction data, we construct an 802K dataset for the instruction tuning stage to facilitate a more equitable comparison. Specifically, we expand the identical 665K instruction data used in LLaVA-1.5 to 802K by sampling an additional 142K instances from publicly accessible datasets. Please refer to the appendix for more details. As shown in Tab. 6, although our model only utilizes 1/25 of the pre-training data compared to VILA, Croc-7B achieves significant performance improvements in 8/10 of the benchmarks. Similarly, Croc-13B also achieves significant performance improvements in 9/10 of the benchmarks. These performance improvements demonstrate that our proposed cross-modal comprehension pretraining stage can facilitate the learning of visual tokens by LLMs, thereby substantially improving the visual comprehension and reasoning capacities of LMMs.
@@ 5 Conclusion
In this paper, we introduce a novel pretraining paradigm for LMMs to enhance the visual comprehension capabilities of LLMs. Our approach incorporates a new cross-modal comprehension stage designed to bridge the gap between the visual and textual domains. Specifically, we develop a dynamically learnable prompt token pool and apply the Hungarian algorithm to replace a portion of the original visual tokens with the most relevant prompt tokens. To further improve the model’s understanding, we propose a mixed attention mechanism that combines bidirectional visual attention with unidirectional textual attention, enabling a more comprehensive interpretation of visual tokens. Additionally, we incorporate a detailed caption generation task, utilizing rich descriptions to improve the LLM’s grasp of visual semantic information. Experimental results show that our method achieves state-of-the-art performance across multiple vision-language benchmarks. We hope that our work offers valuable insights for advancing large multimodal models.
Appendix A Appendix
A.1 Implementation Details
A.1.1 Hyperparameters
@@
A.1.2 Datasets
In the cross-modal comprehension stage, to prevent the degradation of the inherent capabilities of LLM, we also incorporate 300K pure text data. Fig. 8(a) shows the composition of a dataset consisting of text-only data. This data includes 123K, 52K, 30K, 25K, 20K, 25K, 15K, and 10K samples from MathInstruct (Yue et al., 2023), Standford Alpaca (Taori et al., 2023), BELLE (BELLEGroup, 2023), OpenPlatypus (Lee et al., 2023), CodeAlpaca (Chaudhary, 2023), Firefly111https://github.com/yangjianxin1/Firefly, Webqa (Chang et al., 2022), Dolly (Mike et al., 2023).
To facilitate a more equitable comparison with VILA, we construct an 802K instruction dataset categorized into General, OCR, Chart, Math, Text-only, and Other. Fig. 8(b) illustrates the distribution of data across various categories. Specifically, the General dataset comprises 158K, 86K, 83K, 82.5K, 11K, 11K, and 9K samples sourced from LLaVA-Instruct (Liu et al., 2024b), VG (Krishna et al., 2017), VQAv2 (Goyal et al., 2017a), A-OKVQA (Schwenk et al., 2022), ShareGPT4V (Chen et al., 2024b), LAION-GPT4V 222https://huggingface.co/datasets/laion/gpt4v-dataset, and OKVQA (Marino et al., 2019). The OCR dataset includes 80K, 23K, 5K, 3K, and 2K samples from OCRVQA (Mishra et al., 2019), SynDog-EN (Kim et al., 2022), ST-VQA (Biten et al., 2019), IAM (Cheng et al., 2022), and InfoVQA (Mathew et al., 2022). Chart data is drawn from 10K, 18K, and 28K samples taken from DocVQA (Mathew et al., 2021), ChartQA (Masry et al., 2023), and DVQA (Kafle et al., 2018). To enhance the model logical capacity, we respective sample 5K data from Geomverse (Kazemi et al., 2023), Clevr (Johnson et al., 2017), IconQA (Lu et al., 2021b), RAVEN (Huang et al., 2023), Metamath (Yu et al., 2023) and append additional 3K, 1K, and 1K data from Goat (Liu & Low, 2023), Inter-GPS (Lu et al., 2021a), and LIMA (Zhou et al., 2024). Besides, we append 40K pure text from ShareGPT 333https://huggingface.co/datasets/RyokoAI/ShareGPT52K to prevent the degradation of the inherent capabilities of LLM and append 72K, 48K, and 2.5K other data from GQA (Hudson & Manning, 2019) and RefCOCO (Kazemzadeh et al., 2014), AI2D (Kembhavi et al., 2016).
A.2 More Qualitative Results
We present additional qualitative comparisons on visual understanding & reasoning (Fig. 9), image captioning (Fig. 10), and OCR (Fig. 11) tasks, respectively. Benefiting from our proposed cross-modal comprehension stage, the Croc model exhibits enhanced visual comprehension and reasoning capabilities. Meanwhile, it can accurately generate semantic image descriptions and demonstrate robust OCR capabilities.
A.3 Limitation Discussion

### v2

@@ 3.4 Training Pipeline
The overall training pipeline of our Croc model is shown in Fig. 2. Building upon LLaVA-1.5, we propose a two-stage pre-training procedure followed by instruction tuning.
@@
Stage 1.5: Cross-Modal Reconstruction. Building on the cross-modal alignment stage, we introduce the cross-modal reconstruction phase as a pre-training stage. To facilitate comprehensive learning of visual tokens by the LLM, we pretrain the projection layer as well as the LLM. We select 1.2M detailed image-text pairs from the ShareGPT4V [9] dataset, as detailed captions (Fig. 4) enhance the visual token learning of the LLM. To prevent the degradation of the LLM capabilities, we also incorporate 300K pure text data. Please refer to the appendix for more details.
@@ 4 Experiments
Implementation Details. Following LLaVA-1.5, we use CLIP ViT-L/14@336px [56] as the visual encoder and a Vicuna [12] 7B/13B as the LLM. During the cross-modal alignment stage, we set the learning rate for the projection layer to 1e-3. In the cross-modal comprehension stage, we adjust the learning rates to 2e-5 for both the LLM and projection layer and 1.5e-4 for the prompt token pool and projection layer. We set the loss weight to 10, the random mask ratio to , and the size of the prompt token pool to . In the instruction tuning stage, we adopt the learning rate of 2e-5 for LLM and the projection layer. We use AdamW [45] as the optimizer, with a weight decay of and and set to and respectively. We train Croc on NVIDIA A100.
@@
Evaluation Benchmarks. We evaluate Croc across various benchmarks, including 1) General Visual Question Answering: GQA [24], VQAv2 [18], VizWiz [19], LLaVA-Bench (In-the-Wild) [42], MM-Vet [68], RealWorldQA; 2) OCR-Related Question Answering: TextVQA [58], DocVQA [51]; 3) Illusion Benchmarks: POPE [39]; 4) Comprehensive Reasoning Benchmarks: MMBench [44], MME [66], MMStar [10], MM-Vet [68]; 5) Science Visual Question Answering: SciQA [48], AI2D [31].
4.1 Main Results
@@
In Tab. 1, we present a comparison between our model and other methods across various benchmarks. Compared to LLaVA-1.5-7B, Croc-7B demonstrates a 1.9% performance improvement over all benchmarks. Similarly, Croc-13B shows a 1.1% improvement with LLaVA-1.5-13B, thus demonstrating the scalability of a cross-modal reconstruction stage. Next we compare Croc with baselines on instruction-following benchmarks. As shown in Tab. 2, Croc outperforms the other methods. Compared to LLaVA-1.5-7B, Croc-7B exhibits significant performance gains of 1.0%, 15.7, 3.3%, 1.4%, 5.5%, 8.5%, and 4.4% on POPE, MME, MMB, MMB-CN, SEED, LLaVA-W, and MM-Vet. We also evaluate Croc’s performance on multimodal understanding benchmarks in Tab. 3. Croc-7B obtains improvements of 7.4%, 4.4%, 3.8%, and 1.5% over LLaVA-1.5-7B on MMStar, DocVQA, AI2D, and RealWorldQA. Finally, we observe a considerable improvement when increasing the size of the LLM (Croc-13B).
@@
Qualitative Analysis. In the first example shown in Fig. 5, our model captures more detailed information from images, such as specific vehicle models and lower shooting angles. As illustrated in the second example of Fig. 5, Croc can efficiently identifies and describes detailed semantic information, such as the parachute under the aircraft.
@@ 4.2 Ablation Study
Prompt Token Pool Size. The prompt token pool provides a set of jointly learnable visual tokens to replace the masked visual tokens. The tokens in the pool can capture global visual context, thus making the task of reconstruction easier at high masking ratios. This level of difficulty can be further refined using the token pool size. As shown in Fig. 6, we find that the larger the number of tokens in the pool, the easier it is to generate detail captions, resulting in a lower training loss (). We show in Tab. LABEL:table:pool_size, training loss saturates when the pool size is 2,048. In contrast, with pool size of 1, this is functionally equivalent to standard masked generation, which is too difficult and subsequently impeding learning. We observe that increasing the token pool from 1 to 2,048 tokens almost linearly improves performance across all benchmarks. However, expanding the pool to 4,096 tokens, makes the visual token reconstruction too easy, thus degrades the performance.
Visual Token Mask Ratio. The mask ratio of visual tokens directly affects the difficulty of the visual token reconstruction task, and thus significantly affects the effectiveness of pretraining. In Table. LABEL:table:mask_ratio, we report the results of experiments with different mask ratios. Similar to the observations with MAE [20], we find that using a 75% masking ratio yields optimal results in several downstream benchmarks. Lower ratios, e.g., 0%, make the pre-training task too easy, while higher ratios, e.g., 90%, make it too difficult. However, even though a 75% masking ratio is optimal, it must be coupled with an appropriately sized token pool size which is discussed in the previous section.
Nearest Neighbors v.s. Hungarian Matching. To improve the utilization rate of the tokens in the prompt token pool, we use the Hungarian Matching algorithm. This step is important to avoid an over-dependence on a small subset of the token pool, while also enabling a sufficient exploration of the token space. In Fig. 7, we present a comparative analysis of token utilization using both Nearest Neighbors and Hungarian Matching. Due to the Hungarian Matching requirement to select distinct prompt tokens, we observe a significant improvement in the overall utilization of prompt token, leading to a significant improvement across many evaluation benchmarks (see Tab. LABEL:table:matching).
Unidirectional Attention v.s. Mixed Attention. To verify the impact of the mixed attention mechanism, we conduct experiments comparing mixed attention with unidirectional attention used in LLaVA-1.5. As shown in Tab. LABEL:table:attention, our proposed mixed attention mechanism achieves significant performance improvement on all the benchmarks. To explore the influence of different attention mechanisms, we compare the average attention scores of response tokens to visual tokens and language instruction tokens in Fig. 8. We use the same question ”Please describe in detail what is in the picture” for 200 randomly selected images. Croc-7B model exhibits significantly higher attention scores for visual tokens, as indicated by the brighter region in the visual token part of the heatmap, suggesting that it puts more emphasis on understanding and attending to the image content.
Brief v.s. Detailed Caption. To investigate the impact of detailed captions on our proposed cross-modal comprehension stage, we use the BLIP [37] model to generate brief captions for the 1.2M images from the ShareGPT4V [9] dataset. As shown in Tab. LABEL:table:caption, using detailed captions achieves a significant performance improvement. This is mainly because the detailed caption contains rich semantic details of images (Fig. 4), facilitating visual token reconstruction by LLM.
Ablation on Stage, Data, and Objectives. In Tab. LABEL:table:stage1, we show the model’s performance without cross-modal alignment pretraining. The feature misalignment between the LLM and visual encoder makes the visual token reconstruction task more challenging, leading to reduced performance.
In Tab. 5, we present ablation studies on the effects of pretraining data and different training objectives. First, we expand the 558K pretraining dataset of LLaVA-1.5 with an additional 1.5M samples, applying the same training method. This addition improves performance in 4 of 9 datasets, but causes declines in the others, likely due to the altered distribution of the original data.
Next, we add a visual token reconstruction task to enhance the integration of the LLM and visual tokens, improving performance in 5 of 9 datasets.
Further, using detailed caption descriptions improves performance in 8 of 9 datasets. Finally, combining visual token reconstruction with detailed caption generation significantly increases the LLM’s understanding of visual features, reaching the best overall performance.
VILA v.s. Croc.
@@
As VILA incorporates an additional one million instruction data, we construct an 802K dataset for the instruction tuning stage to facilitate a more equitable comparison. Specifically, we expand the identical 665K instruction data used in LLaVA-1.5 to 802K by sampling an additional 142K instances from publicly accessible datasets. Please refer to the appendix for more details. As shown in Tab. 6, although our model only utilizes 1/25 of the pre-training data compared to VILA, Croc-7B achieves significant performance improvements in 8/10 of the benchmarks. Similarly, Croc-13B achieves significant performance improvements in 9/10 of the benchmarks. These performance improvements demonstrate that our proposed cross-modal comprehension pretraining stage can facilitate the learning of visual tokens by LLMs, substantially improving the visual comprehension and reasoning of LMMs.
@@ 5 Conclusion
In this paper, we propose Croc, a pretraining method aimed at improving the visual understanding of LMMs. We introduce a cross-modal comprehension stage that bridges the visual and textual domains, and design a dynamically learnable prompt token pool, which utilizes the Hungarian algorithm to replace some visual tokens with relevant prompts. For better understanding, we combine bidirectional visual attention with unidirectional textual attention. Additionally, we incorporate a caption generation task to improve the model’s grasp of visual semantics. Croc outperforms state-of-the-art by up to across a wide range of evaluation benchmarks and without any architectural modifications.
Supplementary Material
1 Implementation Details
1.1 Hyperparameters
@@
1.2 Datasets
To prevent the degradation of the inherent capabilities of the LLM, we incorporate 300K pure text data during the cross-modal comprehension stage. We show in Fig. 9(a) the composition of the dataset consisting of text-only data. This data includes 124k, 52k, 30k, 25k, 20k, 25k, 15k, 10k samples form MathInstruct [69], Standford Alpaca [59], BELLE [3], OpenPlatypus [35], CodeAlpaca [8], Firefly111https://github.com/yangjianxin1/Firefly, Webqa [6], Dolly [13].
To facilitate a more equitable comparison with VILA, we construct an 800K instruction dataset categorized into General, OCR, Chart, Math, Text-only, and Other. Fig. 9(b) illustrates the distribution of data across various categories. Specifically, the General dataset comprises 158K, 86K, 83K, 46K, 22K, 11K, and 9K samples sourced from LLaVA-Instruct [42], VG [33], VQAv2 [17], ShareGPT4V [9], A-OKVQA [57], LAION-GPT4V 222https://huggingface.co/datasets/laion/gpt4v-dataset, and OKVQA [49]. The OCR dataset includes 80K, 18K, 5K, 3K, and 2K samples from OCRVQA [53], SynDog-EN [32], ST-VQA [4], IAM [11], and InfoVQA [52]. Chart data is drawn from 10K, 18K, and 23K samples taken from DocVQA [51], ChartQA [50], and DVQA [28]. To enhance the model logical capacity, we respective sample 5K data from Geomverse [29], Clevr [27], IconQA [47], RAVEN [23], Metamath [67] and append additional 3K, 1K, 1K data from Goat [43], Inter-GPS [46], and LIMA [70]. Besides, we append 40K pure text from ShareGPT 333https://huggingface.co/datasets/RyokoAI/ShareGPT52K to prevent the degradation of the inherent capabilities of LLM and append 72K, 48K, 2.4K other data from GQA [24] and RefCOCO [30], AI2D [31].
2 More Qualitative Results
We present additional qualitative comparisons on image captioning (Fig. 10), visual understanding & reasoning (Fig. 11), and OCR (Fig. 12) tasks, respectively. Benefiting from our proposed cross-modal comprehension stage, the Croc model exhibits enhanced visual comprehension and reasoning capabilities. Furthermore, it can accurately generate semantic image descriptions and demonstrate robust OCR capabilities.
3 Limitation Discussion